# The Sommerfeld-Rellich Framework for Scattering on Hyperbolic Space: Far-Field Patterns and Inverse Problems

### v2

@@ 4.8. Scattering theory for obstacle problems in hyperbolic space
$u(x)=-\int_{\partial\Omega}\left(\frac{\partial u_{1}(y)}{\partial\nu_{\mathbb{B}^{n}}(y)}G_{-\mu\mathrm{i}}(\rho(x,y))-u_{1}(y)\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\nu_{\mathbb{B}^{n}}(y)}\right)d\sigma_{\mathbb{B}^{n}}(y).$
It is easy to check that $u_{1}$ can be readily extended to an $H^{1}_{loc}(\mathbb{B}^{n}\setminus B_{\mathbb{B}^{n}}(0,R))$ function, which together with the uniqueness of solution of equation (5.2) implies that $u_{1}=u^{s}$. This proves the equivalence of equation (5.2) and equation (4.55). The existence of equation (5.2) follows from standard Fredholm alternative theorem if we choose $B_{R_{0}}$ such that $\mu^{2}$ is not the eigenvalue of the operator $\Delta_{\mathbb{B}^{n}}-\frac{(n-1)^{2}}{4}$ in $B_{\mathbb{B}^{n}}(0,R_{0})\setminus\Omega$ with the Dirichlet boundary on $\partial\Omega$ and Neuman boundary condition on $\partial B_{\mathbb{B}^{n}}(0,R)$. ∎
Through the above lemma, we prove that the total field wave $u$ is existing and unique. Then we accomplish the proof of Theorem 4.6. ∎
@@
We have proved that $u^{s}$ is existing and unique. We claim that $u$ can be written as far-field pattern form. Since $u^{s}$ satisfies hyperbolic radiation condition, applying Green representative formula, for any $x\in\mathbb{B}^{n}\setminus B_{\mathbb{B}^{n}}(0,R)$, we can write
(4.59) $\begin{split}&u^{s}(x)=\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\left(\frac{\partial u^{s}}{\partial\nu_{\mathbb{B}^{n}}(y)}G_{-\mu\mathrm{i}}(\rho(x,y))-\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\nu_{\mathbb{B}^{n}}(y)}u^{s}\right)d\sigma_{\mathbb{B}^{n}}(y)\\ &\ \ +\lim\limits_{t\rightarrow+\infty}\int_{\partial B_{\mathbb{B}^{n}}(0,t)}\left(\frac{\partial u^{s}}{\partial\nu_{\mathbb{B}^{n}}(y)}G_{-\mu\mathrm{i}}(\rho(x,y))-\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\nu_{\mathbb{B}^{n}}(y)}u^{s}\right)d\sigma_{\mathbb{B}^{n}}(y)\\ &=\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\left(\frac{\partial u^{s}}{\partial\rho(y)}G_{-\mu\mathrm{i}}(\rho(x,y))-\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\rho(y)}u^{s}\right)d\sigma_{\mathbb{B}^{n}}(y)\\ &\ \ +\lim\limits_{t\rightarrow+\infty}\int_{\partial B_{\mathbb{B}^{n}}(0,t)}\left(\frac{\partial u^{s}}{\partial\rho(y)}G_{-\mu\mathrm{i}}(\rho(x,y))-\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\rho(y)}u^{s}\right)d\sigma_{\mathbb{B}^{n}}(y).\end{split}$
@@
(4.60) $\begin{split}&\lim\limits_{t\rightarrow+\infty}\int_{\partial B_{\mathbb{B}^{n}}(0,t)}\left(\frac{\partial u^{s}}{\partial\rho(y)}G_{-\mu\mathrm{i}}(\rho(x,y))-\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\rho(y)}u^{s}\right)d\sigma_{\mathbb{B}^{n}}(y)\\ &\ \ =\lim\limits_{t\rightarrow+\infty}\int_{\partial B_{\mathbb{B}^{n}}(0,t)}\left(\left(\frac{\partial u^{s}}{\partial\rho(y)}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})u^{s}\right)G_{-\mu\mathrm{i}}-\left(\frac{\partial G_{-\mu\mathrm{i}}}{\partial\rho(y)}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})G_{-\mu\mathrm{i}}\right)u^{s}\right)d\sigma_{\mathbb{B}^{n}}(y)\\ &\ \ \leq\lim\limits_{t\rightarrow+\infty}\left(\int_{\partial B_{\mathbb{B}^{n}}(0,t)}\left|\left(\frac{\partial u^{s}}{\partial\rho(y)}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})\right)u^{s}\right|^{2}d\sigma_{\mathbb{B}^{n}}(y)\right)^{\frac{1}{2}}\left(\int_{\partial B_{\mathbb{B}^{n}}(0,t)}|G_{-\mu\mathrm{i}}|^{2}d\sigma_{\mathbb{B}^{n}}(y)\right)^{\frac{1}{2}}\\ &\ \ \ \ +\lim\limits_{t\rightarrow+\infty}\left(\int_{\partial B_{\mathbb{B}^{n}}(0,t)}\left|\left(\frac{\partial G_{-\mu\mathrm{i}}}{\partial\rho(y)}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})\right)G_{-\mu\mathrm{i}}\right|^{2}d\sigma_{\mathbb{B}^{n}}(y)\right)^{\frac{1}{2}}\left(\int_{\partial B_{\mathbb{B}^{n}}(0,t)}|u^{s}|^{2}d\sigma_{\mathbb{B}^{n}}(y)\right)^{\frac{1}{2}}\\ &\ \ =0.\end{split}$
@@
$u^{s}(x)=\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\left(\frac{\partial u^{s}}{\partial\rho(y)}G_{-\mu\mathrm{i}}(\rho(x,y))-\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\rho(y)}u^{s}\right)d\sigma_{\mathbb{B}^{n}}(y).$
@@
$\begin{split}&\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\rho(y)}\\ &\ \ =\frac{A_{n,-\mu\mathrm{i}}}{2^{\frac{n-1}{2}}}\frac{\left(\cosh(\frac{\rho(x)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(x)}{2})}\int_{0}^{\pi}(\sin t)^{-2\mu\mathrm{i}}dt\\ &\ \ \ \ \times\frac{\partial}{\partial\rho(y)}\left(\frac{\left(\cosh(\frac{\rho(y)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(y)}{2})}\left(1-2\hat{x}\cdot\tanh(\frac{\rho(y)}{2})+|\tanh(\frac{\rho(y)}{2})|^{2}\right)^{\mu\mathrm{i}-\frac{n-1}{2}}\right)+O\left(\frac{1}{\sinh^{\frac{n+1}{2}}(\rho(x))}\right).\end{split}$
@@
(4.61) $\begin{split}u^{s}&=\frac{\left(\cosh(\frac{\rho(x)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(x)}{2})}\int_{0}^{\pi}(\sin t)^{-2\mu\mathrm{i}}dt\times\frac{A_{n,-\mu\mathrm{i}}}{2^{\frac{n-1}{2}}}\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\frac{\left(\cosh(\frac{\rho(y)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(y)}{2})}\left(1-2\hat{x}\cdot y+|y|^{2}\right)^{\mu\mathrm{i}-\frac{n-1}{2}}\frac{\partial u^{s}}{\partial\rho(y)}d\sigma_{\mathbb{B}^{n}}(y)\\ &\ \ +\frac{A_{n,-\mu\mathrm{i}}}{2^{\frac{n-1}{2}}}\frac{\left(\cosh(\frac{\rho(x)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(x)}{2})}\int_{0}^{\pi}(\sin t)^{-2\mu\mathrm{i}}dt\\ &\ \ \ \ \times\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\frac{\partial}{\partial\rho(y)}\left(\frac{\left(\cosh(\frac{\rho(y)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(y)}{2})}\left(1-2\hat{x}\cdot\tanh(\frac{\rho(y)}{2})+|\tanh(\frac{\rho(y)}{2})|^{2}\right)^{\mu\mathrm{i}-\frac{n-1}{2}}\right)d\sigma_{\mathbb{B}^{n}}(y)\\ &\ \ \ \ +O\left(\frac{1}{\sinh^{\frac{n+1}{2}}(\rho(x))}\right)\\ &=\frac{\left(\cosh(\frac{\rho(x)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(x)}{2})}u_{\infty}^{\mu}(\hat{x},\xi)+O\left(\frac{1}{\sinh^{\frac{n+1}{2}}(\rho(x))}\right),\end{split}$
@@
(4.62) $\begin{split}u_{\infty}^{\mu}(\hat{x},\xi)&=\left(\int_{0}^{\pi}(\sin t)^{-2\mu\mathrm{i}}dt\times\frac{A_{n,-\mu\mathrm{i}}}{2^{\frac{n-1}{2}}}\right)\left(\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\frac{\left(\cosh(\frac{\rho(y)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(y)}{2})}\left(1-2\hat{x}\cdot y+|y|^{2}\right)^{\mu\mathrm{i}-\frac{n-1}{2}}\frac{\partial u^{s}}{\partial\rho(y)}d\sigma_{\mathbb{B}^{n}}(y)\right.\\ &\ \ +\left.\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\frac{\partial}{\partial\rho(y)}\left(\frac{\left(\cosh(\frac{\rho(y)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(y)}{2})}\left(1-2\hat{x}\cdot\tanh(\frac{\rho(y)}{2})+|\tanh(\frac{\rho(y)}{2})|^{2}\right)^{\mu\mathrm{i}-\frac{n-1}{2}}\right)d\sigma_{\mathbb{B}^{n}}(y)\right).\end{split}$
@@ 5.1. Density theorem for Helmholtz equation in hyperbolic space
The eigenfunction $e_{-2\mu,\xi}(x)$ is a smooth solution of the Helmholtz equation in hyperbolic space
$\big(-\Delta_{\mathbb{B}^{n}}-\frac{(n-1)^{2}}{4}\big)v=\mu^{2}v,\ \ x\in\mathbb{B}^{n}.$
For any $g(\xi)\in L^{2}(\mathbb{S}^{n-1})$, it is also easy to check that the function
@@
Obviously, $\widetilde{W_{\mathbb{B}^{n}}^{\mu}}(\Omega)$ is a closed subspace of $C^{\infty}(\Omega)$ with respect to the topology of $C^{\infty}(\Omega)$. We will prove that
@@
[Herglotz approximation theorem in hyperbolic space] $W_{\mathbb{B}^{n}}^{\mu}(\Omega)$ is dense in $\widetilde{W_{\mathbb{B}^{n}}^{\mu}}(\Omega)$ with respect to the topology of $C^{\infty}(\Omega)$.
@@
$\widehat{u^{\prime}}(2\lambda,\xi)=\int_{\mathbb{B}^{n}}u^{\prime}(x)\big(\frac{\sqrt{1-|x|^{2}}}{|x-\xi|}\big)^{n-1-2i\lambda}dV$
@@
$v(x)=D_{n}\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}(\lambda^{2}-\mu^{2})^{-1}\widehat{u^{\prime}}(2\lambda,\xi)\big(\frac{\sqrt{1-|x|^{2}}}{|x-\xi|}\big)^{n-1-2i\lambda}|c(\lambda)|^{-2}d\lambda d\sigma_{\xi}.$
and the $v$ is well defined since $(\lambda^{2}-\mu^{2})^{-1}\widehat{u^{\prime}}(2\lambda,\xi)\in C^{\infty}(\mathbb{R}\times\mathbb{S}^{n-1})$ due to $\widehat{u^{\prime}}(2\mu,\xi)=0$. By the property of Fourier transform in hyperbolic space, we see that $v$ satisfies the Helmholtz equation in hyperbolic space
@@
since $v$ and its derivative are equal to zero on the boundary $\partial\Omega$. This together with the definition of $u$ and equation (5.2) yields that

### v3

@@ 4.8. Scattering theory for obstacle problems in hyperbolic space
$u(x)=-\int_{\partial\Omega}\left(\frac{\partial u_{1}(y)}{\partial\nu_{\mathbb{B}^{n}}(y)}G_{-\mu_{0}\mathrm{i}}(\rho(x,y))-u_{1}(y)\frac{\partial G_{-\mu_{0}\mathrm{i}}(\rho(x,y))}{\partial\nu_{\mathbb{B}^{n}}(y)}\right)d\sigma_{\mathbb{B}^{n}}(y).$
It is easy to check that $u_{1}$ can be readily extended to an $H^{1}_{loc}(\mathbb{B}^{n}\setminus B_{\mathbb{B}^{n}}(0,R))$ function, which together with the uniqueness of solution of equation (4.54) implies that $u_{1}=u_{s}$. This proves the equivalence of equation (4.54) and equation (4.55). The existence of equation (4.54) follows from standard Fredholm alternative theorem if we choose $B_{R_{0}}$ such that $\mu_{0}^{2}$ is not the eigenvalue of the operator $\Delta_{\mathbb{B}^{n}}-\frac{(n-1)^{2}}{4}$ in $B_{\mathbb{B}^{n}}(0,R_{0})\setminus\Omega$ with the Dirichlet boundary on $\partial\Omega$ and Neuman boundary condition on $\partial B_{\mathbb{B}^{n}}(0,R)$. ∎
Through the above lemma, we prove that there exists a unique total field wave $u\in H^{1}_{loc}(\mathbb{B}^{n})$ satisfies equation (4.46). Then we accomplish the proof of Theorem 4.6. ∎
@@
Recall that there exists a unique solution $u_{s}\in H_{loc}^{1}(\mathbb{B}^{n})$ to the scattering system (4.54), we claim that $u$ can be written as far-field pattern form. Since $u^{s}$ satisfies hyperbolic radiation condition, applying Green representative formula, for any $x\in\mathbb{B}^{n}\setminus B_{\mathbb{B}^{n}}(0,R)$, we can write
$\begin{split}&u^{s}=\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\left(\frac{\partial u^{s}}{\partial\nu_{\mathbb{B}^{n}}(y)}G_{-\mu\mathrm{i}}(\rho(x,y))-\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\nu_{\mathbb{B}^{n}}(y)}u^{s}\right)d\sigma_{\mathbb{B}^{n}}(y)\\ &\ \ +\lim\limits_{t\rightarrow+\infty}\int_{\partial B_{\mathbb{B}^{n}}(0,t)}\left(\frac{\partial u^{s}}{\partial\nu_{\mathbb{B}^{n}}(y)}G_{-\mu\mathrm{i}}(\rho(x,y))-\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\nu_{\mathbb{B}^{n}}(y)}u^{s}\right)d\sigma_{\mathbb{B}^{n}}(y)\\ &=\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\left(\frac{\partial u^{s}}{\partial\rho(y)}G_{-\mu\mathrm{i}}(\rho(x,y))-\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\rho(y)}u^{s}\right)d\sigma_{\mathbb{B}^{n}}(y)\\ &\ \ +\lim\limits_{t\rightarrow+\infty}\int_{\partial B_{\mathbb{B}^{n}}(0,t)}\left(\frac{\partial u^{s}}{\partial\rho(y)}G_{-\mu\mathrm{i}}(\rho(x,y))-\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\rho(y)}u^{s}\right)d\sigma_{\mathbb{B}^{n}}(y).\end{split}$
@@
$\begin{split}&\lim\limits_{t\rightarrow+\infty}\int_{\partial B_{\mathbb{B}^{n}}(0,t)}\left(\frac{\partial u^{s}}{\partial\rho(y)}G_{-\mu\mathrm{i}}(\rho(x,y))-\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\rho(y)}u^{s}\right)d\sigma_{\mathbb{B}^{n}}(y)\\ &\ \ =\lim\limits_{t\rightarrow+\infty}\int_{\partial B_{\mathbb{B}^{n}}(0,t)}\left(\left(\frac{\partial u^{s}}{\partial\rho(y)}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})u^{s}\right)G_{-\mu\mathrm{i}}-\left(\frac{\partial G_{-\mu\mathrm{i}}}{\partial\rho(y)}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})G_{-\mu\mathrm{i}}\right)u^{s}\right)d\sigma_{\mathbb{B}^{n}}(y)\\ &\ \ \leq\lim\limits_{t\rightarrow+\infty}\left(\int_{\partial B_{\mathbb{B}^{n}}(0,t)}\left|\left(\frac{\partial u^{s}}{\partial\rho(y)}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})\right)u^{s}\right|^{2}d\sigma_{\mathbb{B}^{n}}(y)\right)^{\frac{1}{2}}\left(\int_{\partial B_{\mathbb{B}^{n}}(0,t)}|G_{-\mu\mathrm{i}}|^{2}d\sigma_{\mathbb{B}^{n}}(y)\right)^{\frac{1}{2}}\\ &\ \ \ \ +\lim\limits_{t\rightarrow+\infty}\left(\int_{\partial B_{\mathbb{B}^{n}}(0,t)}\left|\left(\frac{\partial G_{-\mu\mathrm{i}}}{\partial\rho(y)}-(\mu\mathrm{i}-\frac{n-1}{2})\tanh(\frac{\rho}{2})\right)G_{-\mu\mathrm{i}}\right|^{2}d\sigma_{\mathbb{B}^{n}}(y)\right)^{\frac{1}{2}}\left(\int_{\partial B_{\mathbb{B}^{n}}(0,t)}|u^{s}|^{2}d\sigma_{\mathbb{B}^{n}}(y)\right)^{\frac{1}{2}}\\ &\ \ =0.\end{split}$
@@
$u^{s}=\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\left(\frac{\partial u^{s}}{\partial\rho(y)}G_{-\mu\mathrm{i}}(\rho(x,y))-\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\rho(y)}u^{s}\right)d\sigma_{\mathbb{B}^{n}}(y).$
@@
$\begin{split}&\frac{\partial G_{-\mu\mathrm{i}}(\rho(x,y))}{\partial\rho(y)}\\ &\ \ =\frac{A_{n,-\mu\mathrm{i}}}{2^{\frac{n-1}{2}}}\frac{\left(\cosh(\frac{\rho(x)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(x)}{2})}\int_{0}^{\pi}(\sin t)^{-2\mu\mathrm{i}}dt\\ &\ \ \ \ \times\frac{\partial}{\partial\rho(y)}\left(\frac{\left(\cosh(\frac{\rho(y)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(y)}{2})}\left(1-2\hat{x}\cdot\tanh(\frac{\rho(y)}{2})+|\tanh(\frac{\rho(y)}{2})|^{2}\right)^{\mu\mathrm{i}-\frac{n-1}{2}}\right)+O\left(\frac{1}{\sinh^{\frac{n+1}{2}}(\rho(x))}\right)\end{split}$
@@
$\begin{split}u^{s}&=\frac{\left(\cosh(\frac{\rho(x)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(x)}{2})}\int_{0}^{\pi}(\sin t)^{-2\mu\mathrm{i}}dt\times\frac{A_{n,-\mu\mathrm{i}}}{2^{\frac{n-1}{2}}}\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\frac{\left(\cosh(\frac{\rho(y)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(y)}{2})}\left(1-2\hat{x}\cdot y+|y|^{2}\right)^{\mu\mathrm{i}-\frac{n-1}{2}}\frac{\partial u^{s}}{\partial\rho(y)}d\sigma_{\mathbb{B}^{n}}(y)\\ &\ \ +\frac{A_{n,-\mu\mathrm{i}}}{2^{\frac{n-1}{2}}}\frac{\left(\cosh(\frac{\rho(x)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(x)}{2})}\int_{0}^{\pi}(\sin t)^{-2\mu\mathrm{i}}dt\\ &\ \ \ \ \times\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\frac{\partial}{\partial\rho(y)}\left(\frac{\left(\cosh(\frac{\rho(y)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(y)}{2})}\left(1-2\hat{x}\cdot\tanh(\frac{\rho(y)}{2})+|\tanh(\frac{\rho(y)}{2})|^{2}\right)^{\mu\mathrm{i}-\frac{n-1}{2}}\right)d\sigma_{\mathbb{B}^{n}}(y)\\ &\ \ \ \ +O\left(\frac{1}{\sinh^{\frac{n+1}{2}}(\rho(x))}\right)\\ &=\frac{\left(\cosh(\frac{\rho(x)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(x)}{2})}u_{\infty}^{\mu}(\hat{x},\xi)+O\left(\frac{1}{\sinh^{\frac{n+1}{2}}(\rho(x))}\right),\end{split}$
@@
$\begin{split}u_{\infty}^{\mu}(\hat{x},\xi)&=\left(\int_{0}^{\pi}(\sin t)^{-2\mu\mathrm{i}}dt\times\frac{A_{n,-\mu\mathrm{i}}}{2^{\frac{n-1}{2}}}\right)\left(\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\frac{\left(\cosh(\frac{\rho(y)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(y)}{2})}\left(1-2\hat{x}\cdot y+|y|^{2}\right)^{\mu\mathrm{i}-\frac{n-1}{2}}\frac{\partial u^{s}}{\partial\rho(y)}d\sigma_{\mathbb{B}^{n}}(y)\right.\\ &\ \ +\left.\int_{\partial B_{\mathbb{B}^{n}}(0,R)}\frac{\partial}{\partial\rho(y)}\left(\frac{\left(\cosh(\frac{\rho(y)}{2})\right)^{2\mu\mathrm{i}}}{\cosh^{n-1}(\frac{\rho(y)}{2})}\left(1-2\hat{x}\cdot\tanh(\frac{\rho(y)}{2})+|\tanh(\frac{\rho(y)}{2})|^{2}\right)^{\mu\mathrm{i}-\frac{n-1}{2}}\right)d\sigma_{\mathbb{B}^{n}}(y)\right).\end{split}$
@@
Finally, we note that in Theorems 4.3, 4.5, and 4.6, we considered scattering from compactly supported sources, potentials, and obstacles in the hyperbolic space $\mathbb{B}^{n}$, and established the well-posedness of our scattering theory. It is worth emphasizing that, owing to the local nature of the hyperbolic Sommerfeld radiation condition, the theory can be readily extended to more general asymptotically hyperbolic manifolds (cf. [21, 33]).
@@ 5.1. Density theorem for Helmholtz equation in hyperbolic space
The eigenvalue function $e_{-2\mu_{0},\xi}(x)$ is a smooth solution of the Helmholtz equation in hyperbolic space
$\big(-\Delta_{\mathbb{B}^{n}}-\frac{(n-1)^{2}}{4}\big)v=\mu^{2}v,\ \ x\in\mathbb{B}^{n},$
where $\mu\geq 0$. For any $g(\xi)\in L^{2}(\mathbb{S}^{n-1})$, it is also easy to check that the function
@@
Obviously, $\widetilde{W_{\mathbb{B}^{n}}}^{\mu}(\Omega)$ is a closed subspace of $C^{\infty}(\Omega)$ with respect to the topology of $C^{\infty}(\Omega)$. We prove the Herglotz approximation theorem in hyperbolic space.
@@
$W_{\mathbb{B}^{n}}^{\mu}(\Omega)$ is dense in $\widetilde{W_{\mathbb{B}^{n}}^{\mu}}(\Omega)$ with respect to the topology of $C^{\infty}(\Omega)$.
@@
$\widehat{u^{\prime}}(2\lambda,\xi)=\int_{\mathbb{B}^{n}}u^{\prime}(x)\big(\frac{\sqrt{1-|x|^{2}}}{|x-\xi|}\big)^{n-1-2i\lambda}dV=$
@@
$v(x)=D_{n}\int_{-\infty}^{\infty}\int_{\mathbb{S}^{n-1}}(\lambda^{2}-\mu^{2})^{-1}\widehat{u^{\prime}}(\lambda,\xi)\big(\frac{\sqrt{1-|x|^{2}}}{|x-\xi|}\big)^{n-1-2i\lambda}|c(\lambda)|^{-2}d\lambda d\sigma_{\xi}.$
and the $v$ is well defined since $(\lambda^{2}-\mu^{2})^{-1}\widehat{u^{\prime}}(2\lambda,\xi)\in C^{\infty}(\mathbb{R}\times\mathbb{S}^{n-1})$. By the property of Fourier transform in hyperbolic space, we see that $v$ satisfies the Helmholtz equation in hyperbolic space
@@
since $v$ and its derivative are equal to zero on the boundary $\partial\Omega$. This together with the $u$ satisfying equation
$\big(-\Delta_{\mathbb{B}^{n}}-\frac{(n-1)^{2}}{4}\big)u=\mu^{2}u,\ \ x\in\mathbb{B}^{n}$
yields that